OpenStudy (anonymous):

If y -1 = ln x, what does x equal

OpenStudy (freckles):

use that e^(ln(x))=x , x>0

OpenStudy (anonymous):

so does X= 2.3

OpenStudy (freckles):

Not if y-1=ln(x) solve for using the e^(ln(x))=x for x>0

OpenStudy (freckles):

\[e^{y-1}=e^{\ln(x)}\]